# A Noisy Approach to Intrinsically Mixed-State Topological Order

### v2

@@ V Higher-Form Symmetry and non-unitary exfoliation of Walker-Wang models
Remarkably, this structure exactly parallels that of the vectorized density matrix for an imTO in the doubled Hilbert space, with the ket and bra spaces identified with the top and bottom surfaces of a WW model. Much like the surface states of WW models, the deconfined excitations with support solely on the ket (or bra) space can realize non-modular or chiral TO. The aforementioned weak 1-form symmetries (which act simultaneously on the ket and bra spaces) mirror the effect of the tube-like Wilson surfaces in the WW model, when they terminate on the top and bottom surfaces. Moreover, in the doubled Hilbert space representation of the imTO, the transparent anyons can move freely between the ket and bra spaces, just as the transparent anyons in the WW model can move between the top and bottom surfaces. Indeed, in the doubled Hilbert space, if we condense $(a_{+}a_{-})$, all anyons of the form $(a_{+}^{m}1_{-})$ (for integer $m$) are transparent and are equivalent to anyons of the form $(1_{+}a_{-}^{m})$ via fusion with the condensate. Thus, at least at the level of analogy, decoherence induced imTO provides a physical means of realizing the anomalous surface states of 3D pure state TO as realized by WW models. This lends further credence to our claim that a partial classification of imTOs should be provided by braided tensor categories, as WW models are classified by the same. In this sense, one may view decoherence as a means of “exfoliating” surface states of a 3D TO into a purely 2D mixed-state using only a finite-depth LPQCs. One may imagine applying this non-unitary exfoliation procedure to isolate anomalous surface states of other exotic pure states, an avenue we intend to pursue in future work.
VI Discussion
@@
A pressing issue is to characterize our family of imTOs via their entanglement structure. While the entanglement entropy has previously been studied in mixed-state TO [51, 66], a more natural probe of entanglement in mixed-states is provided by the entanglement negativity which, unlike the entanglement entropy, is a good measure of quantum correlations in a mixed state [67, 68, 69, 70, 71]. In pure state TO, the negativity has been shown to receive universal contributions which are sensitive to the modular data of the TO (namely, the total quantum dimension) [72, 73, 74, 75, 76, 77, 12]. This topological entanglement negativity (TEN) has also been shown to be sensitive to the breakdown of TO in thermal states [78, 12], which the entanglement entropy does not accurately reflect. Since we have shown that imTO is generally characterized by non-modular anyon theories, it is an intriguing question as to what universal data the TEN captures in these states. In one specific instance, Ref. [27] distinguished between the $\mathbb{Z}_{2}^{(0)}$ and $\mathbb{Z}_{2}^{(1)}$ imTOs (obtained via decoherence of the $\mathbb{Z}_{2}$ Toric code) by the respective absence and presence of topological contributions to the negativity. Recalling that these two states correspond to quantum condensates of bosonic and fermionic loops, it is tempting to conjecture that the TEN remains sensitive to the spins of the underlying deconfined anyon excitations. In a forthcoming work, we address more comprehensively the connection between TEN and the braided tensor category structure of imTO. It would likewise be interesting to understand novel decoherence induced negativity transitions [15, 27, 79, 80] that may result from (competition between) the different decoherence channels discussed in this work.
@@
While the entanglement negativity is a good measure of bipartite entanglement, it has recently been understood that pure state TO can be more finely characterized by its tripartite entanglement structure [81, 82, 83, 84, 85, 86, 87, 88]. Specifically, it has been argued that chiral TO supports tripartite entanglement beyond that of the Greenberger-Horne-Zeilinger (GHZ) type [81, 82, 88]. As we have shown, decoherence of the double semion state can induce an imTO characterized by the chiral semion TSSC. Intriguingly, this suggests that decoherence has transmuted one form of many-body tripartite entanglement (i.e. GHZ-like entanglement) into another. It is conceivable that non-unitary processes may stabilize patterns of multipartite entanglement in many-body systems which do not arise naturally in the ground states of local gapped Hamiltonians. Understanding in more detail the multipartite entanglement of imTOs and the transmutation between different classes of entanglement via non-unitary processes promises to be a fruitful direction for further research. In a similar vein, Ref. [89] recently argued that a state that is strongly symmetric with respect to an anomalous 0-form symmetry in 2D must be 4-partite non-separable. It is an intriguing question whether similar constraints exist for systems respecting an anomalous strong 1-form symmetry.
@@
In the spirit of fleshing out the structure of the family of imTOs we have obtained, an important avenue for further development is a more thorough classification of non-Abelian imTO. While we have demonstrated that the process of gauging out via decoherence extends naturally to the non-Abelian case, we do not yet have a comprehensive understanding of the algebraic structure of the resulting imTO, although we have provided compelling evidence that the appropriate mathematical framework is that of braided fusion categories. To that end, it would be prudent to understand more fully the connections with the mixed TQFTs proposed by Zini and Wang [58]. In particular, it remains to be understood whether the class of mixed TQFTs proposed in that work can be realized in a physical setting i.e., by exposing some parent state to local noise. On a related note, it would be interesting to understand whether there exists a non-equilibrium, continuum field theory formulation for describing generic imTO, most likely in the language of the Schwinger-Keldysh path integral.
@@ Acknowledgements.
We are particulary grateful to Dominic Else, Tarun Grover, Tim Hsieh, Kohei Kawabata, Shinsei Ryu, Shengqi Sang, Nati Seiberg, Wilbur Shirley, Nikita Sopenko, Xiao-Qi Sun, Ruben Verresen, and Chong Wang for stimulating discussions. R.S. is especially grateful to Shinsei Ryu for pointing out, and discussions regarding, Ref. [58]. A. P. thanks Fiona Burnell, Sanket Chirame, and Sarang Gopalakrishnan for collaboration on a related project. This material is based upon work supported by a Simons Investigator Grant from the Simons Foundation (Grant No. 566116) awarded to Shinsei Ryu (R. S.), the Sivian Fund at the Institute for Advanced Study, and the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Award Number DE-SC0009988 (A. P.).
Note.—During the completion of this work, we were informed of a forthcoming manuscript [93] which also addresses mixed-state topological order.

### v3

@@ VI Higher-Form Symmetry and non-unitary exfoliation of Walker-Wang models
Remarkably, this structure exactly parallels that of the vectorized density matrix for an imTO in the doubled Hilbert space, with the ket and bra spaces identified with the top and bottom surfaces of a WW model. Much like the surface states of WW models, the deconfined excitations with support solely on the ket (or bra) space can realize non-modular or chiral TO. The aforementioned weak 1-form symmetries (which act simultaneously on the ket and bra spaces) mirror the effect of the tube-like Wilson surfaces in the WW model, when they terminate on the top and bottom surfaces. Moreover, in the doubled Hilbert space representation of the imTO, the transparent anyons can move freely between the ket and bra spaces, just as the transparent anyons in the WW model can move between the top and bottom surfaces. Indeed, in the doubled Hilbert space, if we condense $(a_{+}a_{-})$, all anyons of the form $(a_{+}^{m}1_{-})$ (for integer $m$) are transparent and are equivalent to anyons of the form $(1_{+}a_{-}^{m})$ via fusion with the condensate.
This picture also provides a convenient way of understanding the set of detectable anyons discussed in the preceding section. Indeed, focusing on a single surface of a WW model in a slab geometry, there are two classes of surface excitations which can be detected via braiding with anyons in the surface anyon theory $\mathcal{A}$ – that is to say, detectable via computing the expectation value of a closed Wilson loop of an anyon $d\in\mathcal{A}$. The first are the opaque anyons in $\mathcal{A}$, generated by open Wilson lines on the surface; by definition, the transparent anyons are undetectable via braiding with anyons in $\mathcal{A}$. The second class is generated by open Wilson surfaces in the bulk, the boundaries of which intersect the surface on an open line – the endpoints correspond to point-like excitations on the surface  [^13] While these loop-like excitations are strictly speaking confined as there is an energy cost proportional to the length of the boundary of the Wilson surface, they still nominally exist as excitations in the boundary of the WW model.. These are in one-to-one correspondence with the tube-like Wilson surfaces mentioned above. These two sets of surface excitations exactly correspond to the detectable anyons of the imTO characterized by $\mathcal{A}$ discussed previously. Indeed, in the analogy with the double space picture, the two sets of excitations in the WW model correspond to anyons of the form $(c_{+},c_{-})$ such that $c$ braids non-trivially with some $d\in\mathcal{A}$; the genuine surface Wilson lines in the WW model map to those anyons in the double space such that $c$ braids trivially with the set of gauged out anyons $\hat{\mathcal{A}}$, while the bulk Wilson surfaces generating surface excitations map to those anyons in the double space such that $c$ braids non-trivially with $\hat{\mathcal{A}}$.
Thus, at least at the level of analogy, decoherence induced imTO provides a physical means of realizing anomalous surface states of 3D pure state TO as realized by WW models. This lends further credence to our claim that a partial classification of imTOs is provided by non-modular unitary braided tensor categories, since the classification of WW models includes these. One may view decoherence as a means of “exfoliating” surface states of a 3D TO into a purely 2D mixed-state via finite-depth LPQCs. Indeed, we can also imagine applying such non-unitary exfoliation to isolate anomalous surface states of other exotic pure states, an avenue we intend to pursue in future work.
imTO Phases
Finally, let us comment on the extent to which the imTOs we have discussed thus far constitute genuine mixed-state phases and not simply fixed-point states. We note that this is a subtle question, as there is as of yet no consensus on what constitutes a mixed-state phase of matter. This requires a sharp notion of an equivalence relation that determines when two density matrices lie within the same phase, and there exist several proposals in the literature [14, 21, 72, 67, 73]. For simplicity, we restrict to the Abelian case here. As we have emphasized, the 1-form symmetries (strong and weak) and their associated anomalies encode both the braided fusion category data $\mathcal{A}$ as well as the locally detectable anyon types $\mathcal{L}$. As such, these should constitute invariants of a given mixed-state phase which, in the case of a non-modular $\mathcal{A}$ provides an instance of an imTO phase.
Consider the “two-way connectivity” relation described in Ref. [21]: namely, two short-range correlated mixed-states $\rho_{1}$ and $\rho_{2}$ are in the same mixed-state phase if there exist quasi-local quantum channels $\Sigma_{12}$ and $\Sigma_{21}$ such that $\rho_{1}=\Sigma_{12}\rho_{2}$ and $\rho_{2}=\Sigma_{21}\rho_{1}$ (see [21] for a precise definition of a quasi-local channel). Physically, this relation encodes the fact that while such channels can trivialize long-range correlations, they cannot create them in an arbitrarily short amount of time. Now, note that the $\mathbb{Z}_{2}^{(0)}$ imTO (obtained from $Z$-decoherence of the $\mathbb{Z}_{2}$ Toric code) can be considered to be a purely classical state. Indeed, it is completely separable. Further, when defined on a spatial manifold with trivial genus – the plane or the sphere – one can show this state is two-way connected to a trivial product state and hence, under the equivalence relation of Ref. [21], belongs to the trivial phase. However, as discussed in Section V, there exist local, detectable excitations (i.e. errors) above this state, which reflect the mixed anomaly between the strong electric and weak 1-form symmetries. Moreover, when placed on a torus, the corresponding code space encodes two classical bits of information; while we do not have an explicit proof, this fact suggests that certain states within the code-space cannot be two-way connected to the trivial state via a finite-depth local quantum channel on the torus.
More generally, on physical grounds we expect that logical information (encoded in anomalies between the strong 1-form symmetries) will remain robust under quasi-local quantum channels, since these satisfy a Lieb-Robinson bound and cannot generate arbitrary long-range correlations that destroy this information at infinitesimally small noise rates [57]. On the other hand, classical information (encoded in the mixed strong-weak 1-form anomalies) can be smoothly erased as already seen in the example above. Equivalently, the transparent anyons need not be preserved and, since these are required to capture the complete set of locally detectable anyon types, the set $\mathcal{L}$ is not an invariant under this definition of a mixed-state phase. Note that this subtlety does not arise when $\mathcal{A}$ is modular (since $\mathcal{L}=\mathcal{A}$) and only exists for intrinsically mixed states. Therefore, we contend that a finer equivalence relation is required to accurately describe the invariant data of imTO phases and to distinguish between “classical” imTO phases (such as the $\mathbb{Z}_{2}^{(0)}$ Toric code) from purely trivial states. For instance, an equivalence relation based on the Markov length under local Lindbladian evolution could provide an alternative definition for mixed state phases [67]. We leave a detailed investigation of this matter to future work and speculate no further.
VII Discussion
@@
As we have discussed in Sec. VI, perhaps the most outstanding question is establishing an equivalence relation on the space of short-range correlated mixed-states in 2D that provides a clear notion of imTO phases. Based on our work, we expect that all states within the same phase should share the same 1-form anomaly structure, with imTO phases distinguished by the presence of non-modular strong 1-form symmetries. Importantly, this data encodes not only the (non-modular) braided fusion category $\mathcal{A}$ but also the set of locally detectable anyons $\mathcal{L}$. Since an equivalence relation that requires two-way channel connectivity [21] is only sensitive to the logical information (i.e., anomalies between strong 1-form symmetries), a finer diagnostic is required to also capture the classical information (i.e., mixed strong-weak 1-form anomalies). It is unclear whether requiring two-way channel connectivity on any closed manifold will suffice, or if the equivalence relation based on the Markov length along local Lindblad evolution [67] is more suitable.
@@
A pressing issue is to characterize our family of imTOs via their entanglement structure. While the entanglement entropy has previously been studied in mixed-state TO [58, 74], a more natural probe of entanglement in mixed-states is the entanglement negativity which, unlike the entanglement entropy, is a good measure of quantum correlations in a mixed state [75, 76, 77, 78, 79]. In pure state TO, the negativity receives universal contributions which are sensitive to the modular data of the TO (namely, the total quantum dimension) [80, 81, 82, 83, 84, 85, 12]. This topological entanglement negativity (TEN) has also been shown to be sensitive to the breakdown of TO in thermal states [86, 12], which the entanglement entropy does not accurately reflect. Since we have shown that imTO is generally characterized by non-modular anyon theories, it is an intriguing question as to what universal data the TEN captures in these states. In one specific instance, Ref. [31] distinguished between the $\mathbb{Z}_{2}^{(0)}$ and $\mathbb{Z}_{2}^{(1)}$ imTOs (obtained via decoherence of the $\mathbb{Z}_{2}$ Toric code) by the respective absence and presence of topological contributions to the negativity. Recalling that these two states correspond to quantum condensates of bosonic and fermionic loops, it is tempting to conjecture that the TEN remains sensitive to the spins of the underlying deconfined anyon excitations. In the future, we intend to address more comprehensively the connection between TEN and the braided tensor category structure of imTO. It would likewise be interesting to understand novel decoherence induced negativity transitions [16, 31, 87, 88] that may result from (competition between) the different decoherence channels discussed in this work.
@@
While the entanglement negativity is a good measure of bipartite entanglement, it has recently been understood that pure state TO can be more finely characterized by its tripartite entanglement structure [89, 90, 91, 92, 93, 94, 95, 96]. Specifically, it has been argued that chiral TO supports tripartite entanglement beyond that of the Greenberger-Horne-Zeilinger (GHZ) type [89, 90, 96]. As we have shown, decoherence of the double semion state can induce a mixed state characterized by the chiral semion TO. Intriguingly, this suggests that decoherence has transmuted one form of many-body tripartite entanglement (i.e. GHZ-like entanglement) into another. It is conceivable that non-unitary processes may stabilize patterns of multipartite entanglement in many-body systems which do not arise naturally in the ground states of local gapped Hamiltonians. Understanding in more detail the multipartite entanglement of imTOs and the transmutation between different classes of entanglement via non-unitary processes promises to be a fruitful direction for further research. In a similar vein, Ref. [97] recently argued that a state that is strongly symmetric with respect to an anomalous 0-form symmetry in 2D must be 4-partite non-separable. It is an intriguing question whether similar constraints exist for systems respecting an anomalous strong 1-form symmetry (see Ref. [98] for a recent discussion).
@@
In the spirit of fleshing out the structure of the family of imTOs we have obtained, an important avenue for further development is a more thorough classification of non-Abelian imTO. While we have demonstrated that the process of gauging out via decoherence extends naturally to the non-Abelian case, we do not yet have a comprehensive understanding of the algebraic structure of the resulting imTO, although we have provided compelling evidence that the appropriate mathematical framework is that of non-modular braided fusion categories. To that end, it would be prudent to understand more fully the connections with the mixed TQFTs proposed by Zini and Wang [65]. In particular, it remains to be understood whether the class of mixed TQFTs proposed in that work can be realized in a physical setting i.e., by exposing some parent state to local noise. On a related note, it would be interesting to understand whether there exists a non-equilibrium, continuum field theory formulation for describing generic imTO, most likely in the language of the Schwinger-Keldysh path integral.
@@ Acknowledgements.
We are grateful to Meng Cheng, Tyler Ellison, Dominic Else, Tarun Grover, Tim Hsieh, Kohei Kawabata, Michael Levin, Ruochen Ma, Shinsei Ryu, Shengqi Sang, Nati Seiberg, Wilbur Shirley, Nikita Sopenko, Xiao-Qi Sun, Ruben Verresen, Chong Wang, and Carolyn Zhang for stimulating discussions. R.S. is especially grateful to Shinsei Ryu for pointing out, and discussions regarding, Ref. [65]. A. P. thanks Fiona Burnell, Sanket Chirame, and Sarang Gopalakrishnan for collaboration on related projects. This material is based upon work supported by a Simons Investigator Grant from the Simons Foundation (Grant No. 566116) awarded to Shinsei Ryu (R. S.), the Sivian Fund and the Paul Dirac Fund at the Institute for Advanced Study (A. P.), and the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Award Number DE-SC0009988 (A. P.).
Note.—During the completion of this work, we were informed about Ref. [102] which also addresses mixed-state topological order.